# textwidth=0.8textheight=0.9

### v1

@@ 6.3.2 Convergence Analysis
To understand the impact of network depth on convergence behavior, we compare the training dynamics of different complexities of models in operator learning. Many researchers have studied how network depth influences convergence behavior, such as [75, 56], in the context of KAN and MLP-based models. Pant et al. [75] investigated the convergence behavior of shallow and deep networks by tracking the loss function during training. As shown in Fig. 16(a), when solving the 1D Burgers’ equation, KAN-based models converge faster and reach lower final loss values compared to MLPs. In the shallow setting, MLPs show large loss values, strong oscillations, and early saturation, indicating unstable and inefficient learning. In contrast, shallow KANs not only converge rapidly but also achieve far lower and more stable loss levels. In the deep setting, while both models show reduced oscillations and more stable behavior, KAN still outperforms MLPs in terms of convergence speed and final loss. Abueidda et al. [56] further examine convergence under varying levels of network complexity for both DeepOKAN and DeepONet models. Their study also includes two independent random seeds (seed1 and seed2) for weight initialization, ensuring robustness of the observations. Fig. 16(b) presents the training loss profiles for solving the transient Poisson equation. Across all complexity levels and seeds, DeepOKAN consistently demonstrates lower training loss and faster convergence compared to DeepONet. While DeepONet tends to stabilize early, DeepOKAN maintains superior convergence speed and stability throughout training.
@@
To evaluate how different internal designs of KANs influence model performance, several studies have explored architectural variations and their effects on convergence behavior in operator learning tasks [56, 66, 117]. Toscano et al. [117] investigate multiple extensions of the DeepONet framework by integrating MLPs, cKANs, and KKANs, leading to the development of models like DeepOKKAN. Their study incorporates RBFs within the KKAN structure and further enhances the models using the QR-DeepONet framework [237], which improves training stability and convergence by reparameterizing the trunk network through QR decomposition. As shown in Fig. 16(c), KKANs consistently outperform both cKANs and MLPs in terms of relative error convergence. While cKANs initially converge more rapidly, their performance plateaus over time. Notably, the QR formulation boosts the performance of MLPs and KKANs but has a detrimental effect on cKANs, suggesting that KKANs offer greater robustness and adaptability to optimization strategies. These findings underline the importance of architectural flexibility in achieving stable and efficient training within KAN-based operator learning models.
@@
Integrating KAN into the FNO framework offers another perspective on improving convergence in operator learning. In the EFKAN architecture [115], KAN replaces the MLP used in the traditional EFNO setup. As reported by Wang [115], this modification leads to more stable training and improved convergence behavior. Fig. 16(d) shows that EFNO tends to stop training early, potentially before reaching its optimal solution, while EFKAN continues to refine its predictions and ultimately achieves a lower loss. In addition to its improved learning dynamics, EFKAN also demonstrates faster computational performance compared to conventional numerical solvers like finite difference methods, indicating its potential as an efficient surrogate model for complex geophysical simulations such as magnetotelluric forward modeling.
@@ 7 Challenges and Future Directions
The Kolmogorov–Arnold Network framework provides a mathematically grounded method for approximating high-dimensional functions by combining simpler univariate transformations. While this approach has strong theoretical foundations, its practical implementation presents several open challenges that offer rich opportunities for further research and development. In Fig. 17, we provide an overview of the current landscape of KAN research and highlight key methodological challenges across different model categories. Among the most pressing issues are the high computational cost associated with KANs’ layered structure, the sensitivity of their performance to various hyperparameter choices, and the complexity of their algorithmic design, especially in terms of numerical stability and integration into standard machine learning environments. In what follows, we examine these difficulties in detail and discuss recent methods proposed to improve the efficiency, robustness, and usability of KANs across a range of real-world applications. Moreover, future work can focus on strengthening theoretical foundations, improving performance on complex geometries, and validating KANs on large-scale, real-world engineering and industrial systems to enable more robust and scalable applications.
One significant challenge in implementing the KAN lies in its high computational cost, which arises from the complexity of learning multivariate functions through the composition and addition of univariate functions. This formulation leads to intricate hierarchical structures and adaptive activation mechanisms that require intensive optimization. As a result, training KANs often becomes time-consuming, particularly when applied to high-dimensional problems that demand precise approximations. Fine-grained adjustments and iterative refinements are typically needed to capture the target mappings accurately, further increasing computational overhead. These demands limit the scalability of KANs in large-scale or real-time applications, where computational efficiency is essential. To address these challenges, several recent studies have proposed strategies aimed at reducing both training time and memory usage. For example, the development of separable PIKANs (SPIKANs) [101] has shown that leveraging the separability of physical laws allows for simplified network architectures, resulting in significantly lower computational costs. Another promising direction is domain decomposition [54, 105], where the problem is divided into smaller subdomains, each handled by a reduced-size network. This approach enables parallel or distributed training and decreases the overall complexity of the model, making KANs more suitable for large or distributed systems. While these advances have led to meaningful improvements, achieving broader scalability and faster training remains an active challenge. Continued refinement of architectural principles, training procedures, and hardware-aware implementations will likely play an important role in unlocking the full potential of the Kolmogorov–Arnold framework in applied machine learning contexts.
@@
Another key challenge in KANs lies in their strong sensitivity to hyperparameters, especially those that govern the design and refinement of their compositional architecture. Unlike conventional neural networks, which typically rely on fixed activation functions and standard layer configurations, KANs require precise control over the complexity of univariate basis functions, their adaptive combinations, and the depth of hierarchical compositions. Without well-defined heuristics, setting these hyperparameters often involves a manual and computationally expensive trial-and-error process. Poor choices can result in underfitting due to insufficient representation power or, conversely, lead to excessive training costs due to over-parameterization. In addition, the evolving nature of function approximation within KANs introduces the need for continual adjustment, as the ideal hyperparameter values may shift during training in response to the model’s internal dynamics. To support more efficient and systematic hyperparameter tuning, several strategies rooted in automatic optimization have been proposed. Bayesian optimization techniques [238], for instance, offer a probabilistic approach to exploring the hyperparameter space while minimizing the number of costly model evaluations. These methods are particularly well-suited to KANs, where each training run can be resource-intensive. Similarly, gradient-based hyperparameter optimization frameworks [239] have shown potential for adapting hyperparameters by leveraging gradient signals from the training loss, which could be customized to align with KANs’ compositional structure. Such methods reduce the reliance on exhaustive grid searches and improve efficiency by guiding tuning based on the network’s evolving behavior. By combining these algorithmic tools with a deeper understanding of KAN-specific design principles, researchers are beginning to develop more robust and adaptive approaches to managing complexity in training. Nevertheless, refining these strategies to better match the unique architectural features of KANs remains an area with considerable potential for further refinement and insight.
@@
The third challenge in the practical use of KANs is their lack of built-in support in mainstream deep learning frameworks, such as TensorFlow and PyTorch, which complicates their integration into standard deep learning workflows and limits computational efficiency. The recursive, loop-intensive nature of KANs, stemming from the layered composition of univariate functions, introduces practical difficulties during implementation, particularly when scaling to large models or deploying them in parallel computing environments. These recursive operations often lead to non-trivial memory access patterns and inefficient execution on modern hardware accelerators, which are optimized for regular, feedforward computations. This not only exacerbates challenges related to training time and hyperparameter tuning but also limits the ability to exploit GPU-level parallelism effectively. One promising approach to mitigate this issue involves reformulating recursive expressions using algebraically equivalent, non-recursive structures. For instance, in the case of Chebyshev polynomials, a triangular matrix formulation can replace recursive evaluations, enabling more efficient and parallelizable computations. However, these functions, while theoretically valuable for representing complex nonlinearities, can suffer from numerical instability, particularly when evaluated outside stable intervals or in the presence of noisy gradients. Such issues are especially pronounced in physics-informed applications or differential equation solvers, where accuracy is critical. To improve robustness, researchers have introduced normalization and rescaling techniques that map input variables to compact domains such as [1,1][-1,1][ - 1 , 1 ]. This strategy reduces numerical error and maintains stability during training, especially in models involving trigonometric components or sensitive differential operators [68, 105]. On the implementation side, efforts have also been made to develop more practical and accessible KAN tools. For example, the construction of dedicated KAN modules with optimized computational kernels can streamline training and inference, while minimizing unnecessary overhead. Integration with widely-used libraries through user-friendly application programming interfaces (APIs) and visual debugging tools would further facilitate adoption among practitioners and researchers alike. These enhancements are essential not only for improving usability but also for encouraging more widespread experimentation with KANs in diverse application areas.
@@
The first research direction focuses on developing a rigorous theoretical foundation to better understand the generalization behavior and learning dynamics of KAN-based models. Prior work such as [105] has investigated the training behavior of cPIKANs using the Neural Tangent Kernel framework, providing valuable insights into how architectural choices influence optimization. However, this analysis is limited to a specific case and does not extend to the broader family of PIKAN models or more complex variants like DeepOKANs. A deeper theoretical understanding is required to characterize how the NTK evolves with model complexity and to develop formal connections between the NTK spectrum and prediction error. In addition, NTK-based perspectives represent only one possible lens; other theoretical approaches, such as approximation theory, stability analysis, or generalization error bounds, are essential for a more complete understanding of these models. For example, studies that establish formal generalization bounds for PINNs in inverse PDE settings [240] or derive approximation error estimates for operator networks like FNOs [241] illustrate the type of theoretical analysis that is still lacking for PIKANs and DeepOKANs. In particular, formal characterizations of how model complexity influences generalization, convergence rates, and expressiveness remain largely underdeveloped. Building such theoretical tools could lead to a deeper understanding of KANs and guide the principled design, analysis, and application of these networks across a wide variety of tasks.
@@
The second research direction addresses the challenge of applying KAN-based models to irregular domains and complex geometries that frequently arise in scientific and engineering contexts. The second research direction focuses on enhancing the ability of KAN-based models to operate effectively on irregular domains and complex geometries commonly encountered in scientific and engineering problems. While the Kolmogorov–Arnold structure at the core of PIKANs and DeepOKANs is inherently geometry-agnostic, practical implementations often rely on regular, rectangular grid structures, limiting performance on complex domains. As shown in [53], PIKANs demonstrates limited improvement over traditional MLP-based approaches when applied to irregular geometries, despite its effectiveness on structured ones. A key constraint stems from the rectangular nature of the KAN grid in high dimensions, which aligns more naturally with simple geometries. Moreover, a trade-off exists between capturing geometric detail and maintaining manageable resolution: the more intricate the domain, the finer the grid must be, increasing computational cost and complexity. This tension complicates the balance between geometric fidelity and model simplicity. To improve performance on complex domains, Wang et al. [53] propose several strategies drawn from finite element methods, including h-refinement to locally adapt the grid, isoparametric transformations to map curved domains to simpler reference shapes, and conformal mappings to reparameterize irregular geometries into more manageable ones. While such techniques are promising for PIKANs, handling geometry in DeepOKANs remains relatively underexplored. In particular, although KAN layers are geometry-agnostic, practical deployment in deep operator networks often requires coordinate mappings for trunk inputs or boundary interpolation for the branch network. Embedding richer geometry encodings, such as signed distance functions or shape parameterizations, could help bridge this gap. Furthermore, combining these approaches with domain decomposition or overlapping local expansions may offer a pathway toward more robust and scalable solutions in complex geometrical settings.
@@
The third research direction aims to bridge the gap between current academic implementations of KAN-based models and their deployment in large-scale, real-world engineering and industrial systems. Most current evaluations are conducted on academic-scale PDEs or controlled simulations, often in low-dimensional settings. To transition these models toward real-world deployment, extensive validation on high-performance computing (HPC) platforms is essential, particularly to assess scalability, memory efficiency, parallel training performance, and long-term robustness under real-world conditions. Some early progress in this direction has been made through the introduction of ICKANs [230], which were successfully used to model hyperelastic material behavior directly from measurable data such as strain fields and reaction forces. These models demonstrated strong generalization in finite element simulations involving complex geometries and noise-prone data, confirming the potential of physics-informed KAN-based models in engineering applications. However, broader adoption in large-scale simulations, such as those encountered in digital twin systems [56], design optimization, and uncertainty quantification, will require continued effort in algorithm-hardware alignment, model generality, and physical consistency across diverse scenarios. Advancing this line of research could enable KAN-based models to move beyond academic proof-of-concept and toward robust, scalable tools for industrial use.

### v2

@@ 6.3.2 Convergence Analysis
To understand the impact of network depth on convergence behavior, we compare the training dynamics of different complexities of models in operator learning. Many researchers have studied how network depth influences convergence behavior, such as pant2025mlps , abueidda2025deepokan , in the context of KAN and MLP-based models. Pant et al. pant2025mlps investigated the convergence behavior of shallow and deep networks by tracking the loss function during training. As shown in Fig. 17(a), when solving the 1D Burgers’ equation, KAN-based models converge faster and reach lower final loss values compared to MLPs. In the shallow setting, MLPs show large loss values, strong oscillations, and early saturation, indicating unstable and inefficient learning. In contrast, shallow KANs not only converge rapidly but also achieve far lower and more stable loss levels. In the deep setting, while both models show reduced oscillations and more stable behavior, KAN still outperforms MLPs in terms of convergence speed and final loss. Abueidda et al. abueidda2025deepokan further examine convergence under varying levels of network complexity for both DeepOKAN and DeepONet models. Their study also includes two independent random seeds (seed1 and seed2) for weight initialization, ensuring robustness of the observations. Fig. 17(b) presents the training loss profiles for solving the transient Poisson equation. Across all complexity levels and seeds, DeepOKAN consistently demonstrates lower training loss and faster convergence compared to DeepONet. While DeepONet tends to stabilize early, DeepOKAN maintains superior convergence speed and stability throughout training.
@@
To evaluate how different internal designs of KANs influence model performance, several studies have explored architectural variations and their effects on convergence behavior in operator learning tasks abueidda2025deepokan , shukla2024comprehensive , toscano2025kkans . Toscano et al. toscano2025kkans investigate multiple extensions of the DeepONet framework by integrating MLPs, cKANs, and KKANs, leading to the development of models like DeepOKKAN. Their study incorporates RBFs within the KKAN structure and further enhances the models using the QR-DeepONet framework lee2024training , which improves training stability and convergence by reparameterizing the trunk network through QR decomposition. As shown in Fig. 17(c), KKANs consistently outperform both cKANs and MLPs in terms of relative error convergence. While cKANs initially converge more rapidly, their performance plateaus over time. Notably, the QR formulation boosts the performance of MLPs and KKANs but has a detrimental effect on cKANs, suggesting that KKANs offer greater robustness and adaptability to optimization strategies. These findings underline the importance of architectural flexibility in achieving stable and efficient training within KAN-based operator learning models.
@@
Integrating KAN into the FNO framework offers another perspective on improving convergence in operator learning. In the EFKAN architecture wang2025efkan , KAN replaces the MLP used in the traditional EFNO setup. As reported by Wang wang2025efkan , this modification leads to more stable training and improved convergence behavior. Fig. 17(d) shows that EFNO tends to stop training early, potentially before reaching its optimal solution, while EFKAN continues to refine its predictions and ultimately achieves a lower loss. In addition to its improved learning dynamics, EFKAN also demonstrates faster computational performance compared to conventional numerical solvers like finite difference methods, indicating its potential as an efficient surrogate model for complex geophysical simulations such as magnetotelluric forward modeling.
@@ 7 Challenges and Future Directions
While the Kolmogorov–Arnold Network frameworks offer a mathematically grounded approach for approximating high-dimensional functions through compositions of simple univariate transformations, several challenges still hinder its broader adoption. These include scalability limitations due to computational cost, interpretability issues, sensitivity to hyperparameters, and difficulties in applying KANs to complex physical systems. In this section, we highlight key challenges and discuss promising directions for future research. These include strengthening the theoretical foundations, improving performance on complex geometries, and validating KANs on large-scale, real-world engineering and industrial applications.
7.1 Methodological Challenges
It is essential to acknowledge the limitations of KANs when compared with MLPs. As comprehensively reviewed by Ji et al. ji2024comprehensive , several studies have shown that the performance of KANs is not universally superior. Zeng et al. zeng2024kan demonstrated that while KANs may exhibit faster convergence and improved fitting on certain irregular or noisy functions, they tend to underperform relative to MLPs for functions with jump discontinuities or non-differentiable points. Their results also indicate that KANs are more sensitive to noise, with test errors increasing noticeably under higher noise levels. KANs generally require significantly greater computational cost and hardware resources, and in some tasks, the training time can exceed that of MLPs by up to two orders of magnitude. In addition, Pourkamali-Anaraki pourkamali2024kolmogorov reported that KANs experience substantial degradation in low-data regimes, particularly when data are scarce or unevenly distributed. Fig. 18 provides an overview of the current landscape of KAN research and highlight key methodological challenges across different model categories. Among the most pressing issues are the high computational cost associated with KANs’ layered structure, the sensitivity of their performance to various hyperparameter choices, and the complexity of their algorithmic design, especially in terms of numerical stability and integration into standard machine learning environments. Table 8 also presents a comparative summary of seminal KAN developments in terms of these four main methodological challenges across data-driven, physics-informed, and deep-operator learning paradigms. In what follows, we examine these challenges in detail and discuss recent approaches proposed to improve the efficiency, robustness, and usability of KANs across a range of real-world applications.
@@ 7.1 Methodological Challenges
Table 8: Comparative summary of major KAN developments in the context of methodological challenges across data-driven, physics-informed, and deep-operator learning paradigms.
Model Name Accuracy Convergence Theoretical Analysis Computational Analysis Code Ref
(I) Data-driven
KAN $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ liu2024kan
Chebyshev KAN $\checkmark$ $\checkmark$ $\times$ $\times$ ss2024chebyshev
FastKAN (RBF) $\checkmark$ $\times$ $\times$ $\checkmark$ li2024kolmogorov
Wav-KAN $\checkmark$ $\times$ $\times$ $\times$ bozorgasl2405wav
(II) Physics-informed
PIKAN $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ wang2024kolmogorov
SPIKAN $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ jacob2024spikans
Scaled-cPIKAN $\checkmark$ $\checkmark$ $\checkmark$ $\times$ mostajeran2025scaled
KKANs $\checkmark$ $\checkmark$ $\times$ $\checkmark$ toscano2025kkans
(III) Deep-operators
DeepOKAN $\checkmark$ $\checkmark$ $\times$ $\times$ abueidda2025deepokan
Hybrid Operators $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ kiyani2025predicting
One significant challenge in implementing the KAN lies in its high computational cost, which arises from the complexity of learning multivariate functions through the composition and addition of univariate functions. This formulation leads to intricate hierarchical structures and adaptive activation mechanisms that require intensive optimization. As a result, training KANs often becomes time-consuming, particularly when applied to high-dimensional problems that demand precise approximations. Fine-grained adjustments and iterative refinements are typically needed to capture the target mappings accurately, further increasing computational overhead. These demands limit the scalability of KANs in large-scale or real-time applications, where computational efficiency is essential. To address these challenges, several recent studies have proposed strategies aimed at reducing both training time and memory usage. For example, the development of separable PIKANs (SPIKANs) jacob2024spikans has shown that leveraging the separability of physical laws allows for simplified network architectures, resulting in significantly lower computational costs. Another promising direction is domain decomposition howard2024finite , faroughi2025neural , where the problem is divided into smaller subdomains, each handled by a reduced-size network. This approach enables parallel or distributed training and decreases the overall complexity of the model, making KANs more suitable for large or distributed systems. Moreover, within deep-operator frameworks, Wu & Lin wu2025po introduced the Chunk-rational KAN (CKAN) and integrated it into the Physics-informed deep-operator architecture to enhance scalability and enforce physical consistency in PDE-based learning. Their study demonstrated that conventional KANs exhibit quadratic parameter growth and extremely high floating-point operations (FLOPs). By incorporating chunk-wise parameter sharing, CKAN achieves a reduction in FLOPs by approximately three orders of magnitude while preserving parameter complexity comparable to MLP baselines and maintaining, or even improving, predictive accuracy, with only minor increases in inference latency. While these advances have led to meaningful improvements, achieving broader scalability and faster training remains an active challenge. Continued refinement of architectural principles, training procedures, and hardware-aware implementations will likely play an important role in unlocking the full potential of the Kolmogorov–Arnold framework in applied machine learning contexts.
@@
Another key challenge in KANs lies in their strong sensitivity to hyperparameters, especially those that govern the design and refinement of their compositional architecture. Unlike conventional neural networks, which typically rely on fixed activation functions and standard layer configurations, KANs require precise control over the complexity of univariate basis functions, their adaptive combinations, and the depth of hierarchical compositions. Without well-defined heuristics, setting these hyperparameters often involves a manual and computationally expensive trial-and-error process. Poor choices can result in underfitting due to insufficient representation power or, conversely, lead to excessive training costs due to over-parameterization. In addition, the evolving nature of function approximation within KANs introduces the need for continual adjustment, as the ideal hyperparameter values may shift during training in response to the model’s internal dynamics. To support more efficient and systematic hyperparameter tuning, several strategies rooted in automatic optimization have been proposed. Bayesian optimization techniques snoek2012practical , for instance, offer a probabilistic approach to exploring the hyperparameter space while minimizing the number of costly model evaluations. These methods are particularly well-suited to KANs, where each training run can be resource-intensive. Similarly, gradient-based hyperparameter optimization frameworks lorraine2020optimizing have shown potential for adapting hyperparameters by leveraging gradient signals from the training loss, which could be customized to align with KANs’ compositional structure. Such methods reduce the reliance on exhaustive grid searches and improve efficiency by guiding tuning based on the network’s evolving behavior. By combining these algorithmic tools with a deeper understanding of KAN-specific design principles, researchers are beginning to develop more robust and adaptive approaches to managing complexity in training. Nevertheless, refining these strategies to better match the unique architectural features of KANs remains an area with considerable potential for further refinement and insight.
@@
The third challenge in the practical use of KANs is their lack of built-in support in mainstream deep learning frameworks, such as TensorFlow and PyTorch, which complicates their integration into standard deep learning workflows and limits computational efficiency. The recursive, loop-intensive nature of KANs, stemming from the layered composition of univariate functions, introduces practical difficulties during implementation, particularly when scaling to large models or deploying them in parallel computing environments. These recursive operations often lead to non-trivial memory access patterns and inefficient execution on modern hardware accelerators, which are optimized for regular, feedforward computations. This not only exacerbates challenges related to training time and hyperparameter tuning but also limits the ability to exploit GPU-level parallelism effectively. One promising approach to mitigate this issue involves reformulating recursive expressions using algebraically equivalent, non-recursive structures. For instance, in the case of Chebyshev polynomials, a triangular matrix formulation can replace recursive evaluations, enabling more efficient and parallelizable computations. However, these functions, while theoretically valuable for representing complex nonlinearities, can suffer from numerical instability, particularly when evaluated outside stable intervals or in the presence of noisy gradients. Such issues are especially pronounced in physics-informed applications or differential equation solvers, where accuracy is critical. To improve robustness, researchers have introduced normalization and rescaling techniques that map input variables to compact domains such as [1,1][-1,1]. This strategy reduces numerical error and maintains stability during training, especially in models involving trigonometric components or sensitive differential operators mostajeran2025scaled , faroughi2025neural . On the implementation side, efforts have also been made to develop more practical and accessible KAN tools. For example, the construction of dedicated KAN modules with optimized computational kernels can streamline training and inference, while minimizing unnecessary overhead. Integration with widely-used libraries through user-friendly application programming interfaces (APIs) and visual debugging tools would further facilitate adoption among practitioners and researchers alike. These enhancements are essential not only for improving usability but also for encouraging more widespread experimentation with KANs in diverse application areas.
7.2 Future Direction
@@
The first research direction focuses on developing a rigorous theoretical foundation to better understand the generalization behavior and learning dynamics of KAN-based models. Studies investigating the feature-learning properties of KANs, mainly through NTK analysis, remain limited and are mostly confined to specific cases mostajeran2025scaled , faroughi2025neural , rigas2025initialization . Prior work such as faroughi2025neural has investigated the training behavior of cPIKANs using the Neural Tangent Kernel framework, providing valuable insights into how architectural choices influence optimization. However, this analysis is limited to a specific case and does not extend to the broader family of PIKAN models or more complex variants like DeepOKANs.A deeper theoretical understanding is required to characterize how the NTK evolves with model complexity and to develop formal connections between the NTK spectrum and prediction error. In addition, NTK-based perspectives represent only one possible lens; other theoretical approaches, such as approximation theory, stability analysis, or generalization error bounds, are essential for a more complete understanding of these models. Beyond NTK-based analyses, some studies have explored the theoretical foundations of practical KANs from an approximation-theoretic perspective, establishing optimal approximation guarantees kratsios2025kolmogorov , explicit error bounds warin2024p1 , and constructive results that mitigate the curse of dimensionality for various spline- and polynomial-based architectures fakhoury2025expressivity . Similarly, studies that establish formal generalization bounds for PINNs in inverse PDE settings mishra2022estimates or derive approximation error estimates for operator networks like FNOs kovachki2021universal illustrate the type of theoretical analysis that is still lacking for PIKANs and DeepOKANs. In particular, formal characterizations of how model complexity influences generalization, convergence rates, and expressiveness remain largely underdeveloped. Building such theoretical tools could lead to a deeper understanding of KANs and guide the principled design, analysis, and application of these networks across a wide variety of tasks.
@@
The second research direction addresses the challenge of applying KAN-based models to irregular domains and complex geometries that frequently arise in scientific and engineering contexts. The second research direction focuses on enhancing the ability of KAN-based models to operate effectively on irregular domains and complex geometries commonly encountered in scientific and engineering problems. While the Kolmogorov–Arnold structure at the core of PIKANs and DeepOKANs is inherently geometry-agnostic, practical implementations often rely on regular, rectangular grid structures, limiting performance on complex domains. As shown in wang2024kolmogorov , PIKANs demonstrates limited improvement over traditional MLP-based approaches when applied to irregular geometries, despite its effectiveness on structured ones. A key constraint stems from the rectangular nature of the KAN grid in high dimensions, which aligns more naturally with simple geometries. Moreover, a trade-off exists between capturing geometric detail and maintaining manageable resolution: the more intricate the domain, the finer the grid must be, increasing computational cost and complexity. This tension complicates the balance between geometric fidelity and model simplicity. To improve performance on complex domains, Wang et al. wang2024kolmogorov propose several strategies drawn from finite element methods, including h-refinement to locally adapt the grid, isoparametric transformations to map curved domains to simpler reference shapes, and conformal mappings to reparameterize irregular geometries into more manageable ones. While such techniques are promising for PIKANs, handling geometry in DeepOKANs remains relatively underexplored. In particular, although KAN layers are geometry-agnostic, practical deployment in deep-operator networks often requires coordinate mappings for trunk inputs or boundary interpolation for the branch network. Embedding richer geometry encodings, such as signed distance functions or shape parameterizations, could help bridge this gap. Furthermore, combining these approaches with domain decomposition or overlapping local expansions may offer a pathway toward more robust and scalable solutions in complex geometrical settings.
@@
The third research direction aims to bridge the gap between current academic implementations of KAN-based models and their deployment in large-scale, real-world engineering and industrial systems. Most current evaluations are conducted on academic-scale PDEs or controlled simulations, often in low-dimensional settings. To transition these models toward real-world deployment, extensive validation on high-performance computing (HPC) platforms is essential, particularly to assess scalability, memory efficiency, parallel training performance, and long-term robustness under real-world conditions. Some early progress in this direction has been made through the introduction of ICKANs thakolkaran2025can , which were successfully used to model hyperelastic material behavior directly from measurable data such as strain fields and reaction forces. These models demonstrated strong generalization in finite element simulations involving complex geometries and noise-prone data, confirming the potential of physics-informed KAN-based models in engineering applications. However, broader adoption in large-scale simulations, such as those encountered in digital twin systems abueidda2025deepokan , design optimization, and uncertainty quantification, will require continued effort in algorithm-hardware alignment, model generality, and physical consistency across diverse scenarios. Advancing this line of research could enable KAN-based models to move beyond academic proof-of-concept and toward robust, scalable tools for industrial use.